# ZeroSearch: Incentivize the Search Capability of LLMs without Searching

### v1

@@ Abstract
Effective information searching is essential for enhancing the reasoning and generation capabilities of large language models (LLMs). Recent research has explored using reinforcement learning (RL) to improve LLMs’ search capabilities by interacting with live search engines in real-world environments. While these approaches show promising results, they face two major challenges: (1) Uncontrolled Document Quality: The quality of documents returned by search engines is often unpredictable, introducing noise and instability into the training process. (2) Prohibitively High API Costs: RL training requires frequent rollouts, potentially involving hundreds of thousands of search requests, which incur substantial API expenses and severely constrain scalability. To address these challenges, we introduce ZeroSearch, a reinforcement learning framework that incentivizes the search capabilities of LLMs without interacting with real search engines. Our approach begins with lightweight supervised fine-tuning to transform the LLM into a retrieval module capable of generating both relevant and noisy documents in response to a query. During RL training, we employ a curriculum-based rollout strategy that incrementally degrades the quality of generated documents, progressively eliciting the model’s reasoning ability by exposing it to increasingly challenging retrieval scenarios. Extensive experiments demonstrate that ZeroSearch effectively incentivizes the search capabilities of LLMs using a 3B LLM as the retrieval module. Remarkably, a 7B retrieval module achieves comparable performance to the real search engine, while a 14B retrieval module even surpasses it. Furthermore, it generalizes well across both base and instruction-tuned models of various parameter sizes and is compatible with a wide range of RL algorithms.
@@ 1 Introduction
Recently, reinforcement learning (RL) has emerged as a promising strategy for further improving LLM performance by enhancing their reasoning and decision-making capabilities [7, 9]. Notably, RL-based models such as OpenAI-o1 and DeepSeek-R1 have demonstrated substantial gains in logical inference and iterative reasoning—achieved purely through reward-driven learning, without relying on explicit step-by-step supervision [20]. Within this paradigm, several studies have explored using RL to train policy models that can more effectively search for relevant information. Representative examples include Search-R1 [17], R1-Searcher [35], and ReSearch [19]. Notably, DeepResearcher [47] introduces live interaction with commercial search engines such as Google, allowing models to train in an environment that closely mirrors real-world web search. Despite these advancements, integrating RL with real-world search scenarios presents significant challenges: (1) Uncontrolled Document Quality: The quality of documents retrieved from live search engines is often unpredictable, introducing noise and instability into the training process. (2) Prohibitively High API Costs: RL training requires frequent rollouts, potentially involving hundreds of thousands of API calls, which incur substantial financial costs and severely limit scalability.
@@
To address these challenges, we propose ZeroSearch—a reinforcement learning framework that enables LLMs to learn search strategies without interacting with real search engines. Our key insight is that LLMs have acquired extensive world knowledge during large-scale pretraining and are capable of generating relevant documents given a search query [43]. The primary difference between a real search engine and a simulation LLM lies in the textual style of the returned content. However, with lightweight supervised fine-tuning, even relatively small LLMs can effectively simulate the behavior of real search engines. In addition to eliminating API costs, an important advantage of using LLMs for document generation is the ability to control document quality. Specifically, during supervised fine-tuning, documents that lead to correct or incorrect answers are distinguished through prompt design, enabling the simulation LLM to learn to generate either relevant or noisy documents simply by adjusting a few words in the prompt. Building on this, we introduce a curriculum rollout mechanism during training, in which the quality of the generated documents is gradually degraded over time to simulate increasingly challenging retrieval scenarios. This allows the policy model to first learn basic output formats and task requirements before progressively adapting to more challenging and noisy retrieval scenarios. More importantly, ZeroSearch exhibits strong scalability: increasing the number of GPUs significantly accelerates the generation throughput of the simulation LLM, thereby enabling efficient large-scale rollout. Empirical results show that even a 3B LLM used as the simulated search engine can effectively incentivize the policy model’s search capabilities. A 7B retrieval module achieves performance comparable to Google Search, while a 14B retrieval module even surpasses it. ZeroSearch is compatible with both base and instruction-tuned models of various parameter sizes, removing the need for separate supervised warm-up stages. Moreover, it integrates seamlessly with widely used RL algorithms, including Proximal Policy Optimization (PPO) [31], Group Relative Policy Optimization (GRPO) [32, 7], and Reinforce++ [10].
@@
  • We propose ZeroSearch, a novel reinforcement learning framework that incentivizes the search capability of LLMs without interacting with real search engines.
  • Through supervised fine-tuning, we transform the LLM into a retrieval module capable of generating both relevant and noisy documents in response to a query. We further introduce a curriculum rollout mechanism to progressively elicit the model’s reasoning ability by exposing it to increasingly challenging retrieval scenarios.
@@ 2.2 Learning to Search through Reinforcement Learning
Recently, reinforcement learning (RL) has emerged as a promising paradigm for enhancing the reasoning capabilities of LLMs [7, 9]. Notable RL-based models such as OpenAI-o1 and DeepSeek-R1 have demonstrated remarkable capabilities in logical inference and iterative reasoning, purely driven by reward signals without explicit step-by-step supervision [20]. Several studies have also explored RL techniques specifically designed to train models for effective information retrieval. For instance, Search-R1 [17] employs reinforcement learning to autonomously generate multiple search queries during step-by-step reasoning. Similarly, R1-Searcher [35] proposes a two-stage, outcome-based RL method aimed at enhancing search capabilities. ReSearch [19] leverages RL to teach models to reason through searches, entirely without supervision on intermediate reasoning steps. However, these methods usually employ static, local textual corpora such as Wikipedia and fail to capture the complexities of real-world interaction. To bridge this gap, DeepResearcher [47] introduces direct interaction with commercial search engines such as Google, allowing training environments that closely approximate real-world search scenarios. While achieving superior performance, these real-time retrieval methods face significant challenges, including unpredictable document quality, prohibitive high API costs that adversely affect system scalability. To address these limitations, we propose ZeroSearch, a method leveraging an LLM to simulate real-time search, effectively removing dependence on costly, rate-limited real search APIs. Through lightweight supervised fine-tuning, ZeroSearch allows explicit control over document quality and implements a curriculum rollout mechanism that enhances training stability and robustness.
@@ 3.3 Search Simulation Tuning
To bridge this gap, we propose a lightweight supervised fine-tuning (SFT) procedure. Specifically, we first collect interaction trajectories by prompting the LLM to engage with a real search engine in a multi-turn manner until a final answer is reached. Trajectories that yield correct answers are labeled as positive, indicating that the retrieved documents were useful, while those leading to incorrect answers are labeled as negative, indicating noisy retrievals.
Then, we extract query-document pairs from both positive and negative trajectories and perform lightweight SFT to improve the LLM’s ability to simulate real search engines. As shown in Table 2, the distinction between useful and noisy retrievals is achieved by adjusting a few words in the prompt. Besides, we also incorporate the input question and its corresponding answer into the prompt to broaden the knowledge boundary of LLMs. After fine-tuning, the LLM is capable of generating both useful and noisy documents, enabling dynamic document quality control during rollout.
@@ 3.6 Training Algorithm
Our approach is compatible with a wide range of reinforcement learning algorithms, including Proximal Policy Optimization (PPO) [31], Group Relative Policy Optimization (GRPO) [32, 7], and Reinforce++ [10], each offering distinct advantages for optimizing retrieval-augmented reasoning.
In ZeroSearch, the rollout sequence comprises both tokens generated by the policy model and document tokens returned by the simulation LLM. Applying the same optimization procedure uniformly across both types of tokens can lead to training instability, as the retrieved content is externally generated and not directly controlled by the policy model.
To mitigate this, we introduce a loss masking mechanism for retrieved tokens, ensuring that gradients are only computed with respect to the model’s own outputs. This strategy stabilizes the RL training process while preserving the effectiveness of retrieval-augmented generation.
@@ 4.3 Experimental Setup
We conduct experiments using three model families: Qwen-2.5-7B (Base/Instruct) and Qwen-2.5-3B (Base/Instruct) [40], as well as LLaMA-3.2-3B (Base/Instruct) [4]. To simulate real-world retrieval scenarios, we utilize Google Web Search via the SerpAPI [^1] https://serpapi.com/ as the external search engine. The number of retrieved documents is fixed at five across all methods to ensure a fair comparison.
@@
To train the simulation LLM, we conduct a lightweight SFT using Qwen-2.5-3B, Qwen-2.5-7B, and Qwen-2.5-14B as the backbones. The learning rate is set to be 1e-6. To train ZeroSearch, we adopt two reinforcement learning algorithms: GRPO and PPO. In the GRPO setting, the policy LLM is trained with a learning rate of 1e-6, and five responses are sampled per prompt. In the PPO setting, the policy LLM is trained with a learning rate of 1e-6, while the value model is trained with a separate learning rate of 1e-5. We apply Generalized Advantage Estimation (GAE) with hyperparameters $\lambda=1$ and $\gamma=1$. Unless otherwise specified, GRPO is used as the default reinforcement learning algorithm, and Qwen-2.5-14B is used as the default simulation LLM throughout all experiments.
@@ 4.4 Performance
R1-base 25.15 43.18 22.29 21.02 28.46 9.76 24.80 24.95
R1-instruct 25.25 42.68 27.81 20.45 26.83 8.33 27.05 25.49
Search-R1-base 41.51 60.53 51.02 32.25 36.31 16.39 28.00 38.00
Search-R1-inst 41.46 62.17 49.80 34.55 34.22 19.43 33.06 39.24
@@
R1-base 21.46 34.02 16.29 22.67 28.57 5.51 7.20 19.39
R1-instruct 17.41 33.13 18.40 18.90 25.66 4.69 15.32 19.07
Search-R1-base 41.13 61.22 40.73 27.64 31.97 12.17 12.40 32.47
Search-R1-inst 40.82 59.11 42.76 30.83 31.10 8.37 13.01 32.28
@@
R1-base 26.63 42.31 20.16 22.56 28.31 7.36 15.45 23.25
R1-instruct 32.32 52.43 25.98 19.51 29.21 10.34 21.95 27.39
Search-R1-base 40.85 61.68 44.47 28.25 32.66 13.74 12.10 33.39
Search-R1-inst 43.72 56.44 45.49 22.45 23.37 9.27 18.70 31.35
@@ 5.1 Compare ZeroSearch with Real Search Engine
We compare the reward curves of ZeroSearch and Search-R1 (using a real search engine) on LLaMA-3.2-3B, as shown in Figures 2(a) and 2(b). Several key observations can be made:
@@
ZeroSearch achieves a more stable and smoother learning curve. As shown in Figure 2(b), ZeroSearch initially lags behind Search-R1 but eventually surpasses it with much less fluctuation, thanks to the curriculum mechanism that helps the model gradually master search tool usage.
@@
(a) LLaMA-3.2-3B-Base
(b) LLaMA-3.2-3B-Inst
@@
Figure 2: (a-b): Reward curve comparison between ZeroSearch and Search-R1 using LLaMA-3.2-3B. (c): Interaction turns and reward progression during training of LLaMA-3.2-3B-base.
@@ 5.2 Choice of Simulation LLMs
Table 4: Performance of simulated search engines using different LLM configurations. We compare prompt-based and fine-tuned simulation LLMs (3B to 14B) against Google Search.
In this section, we investigate how different simulation engine configurations affect performance, including prompt-based and fine-tuned LLMs ranging from 3B to 14B parameters. Based on the results in Table 4, we make the following observations:
@@
Second, fine-tuned simulation engines significantly outperform prompt-based ones. Although prompt-based methods are explicitly guided to mimic the response style of a real search engine, a substantial distribution gap remains, leading to inferior performance.
Third, performance improves consistently with increasing model size. Larger simulation LLMs not only exhibit stronger simulation capabilities but also more accurately distinguish between relevant and irrelevant documents, thereby enabling more effective curriculum learning during training.
@@ 5.3 Interaction Turns Study
In this section, we analyze the training dynamics of ZeroSearch by examining both reward progression and the number of interaction turns throughout the training, using the LLaMA3.2-3B-Base model. The results are shown in Figure 2(c).
@@
As training progresses, both the number of interaction turns and the reward curve increase sharply and then stabilize. This is primarily because the policy model becomes capable of effectively retrieving relevant documents and ultimately achieving correct answers, resulting in higher rewards. Notably, although the reward appears stable in the later stages of training, the underlying task difficulty continues to rise due to the curriculum mechanism. Consequently, the policy must continuously refine its strategy and improve its reasoning ability to maintain consistent performance.
5.4 Different RL algorithms: PPO vs. GRPO

### v2

@@ Abstract
Effective information searching is essential for enhancing the reasoning and generation capabilities of large language models (LLMs). Recent research has explored using reinforcement learning (RL) to improve LLMs’ search capabilities by interacting with live search engines in real-world environments. While these approaches show promising results, they face two major challenges: (1) Uncontrolled Document Quality: The quality of documents returned by search engines is often unpredictable, introducing noise and instability into the training process. (2) Prohibitively High API Costs: RL training requires frequent rollouts, potentially involving hundreds of thousands of search requests, which incur substantial API expenses and severely constrain scalability. To address these challenges, we introduce ZeroSearch, a novel RL framework that incentivizes the capabilities of LLMs to use a real search engine with simulated searches during training. Our approach begins with lightweight supervised fine-tuning to transform the LLM into a retrieval module capable of generating both useful and noisy documents in response to a query. During RL training, we employ a curriculum-based rollout strategy that incrementally degrades the quality of generated documents, progressively eliciting the model’s reasoning ability by exposing it to increasingly challenging retrieval scenarios. Extensive experiments demonstrate that ZeroSearch effectively incentivizes the search capabilities of LLMs using a 3B LLM as the retrieval module. Remarkably, a 7B retrieval module achieves comparable performance to the real search engine, while a 14B retrieval module even surpasses it. Furthermore, it generalizes well across both base and instruction-tuned models of various parameter sizes and is compatible with a wide range of RL algorithms.
@@ 1 Introduction
Recently, reinforcement learning (RL) has emerged as a promising strategy for further improving LLM performance by enhancing their reasoning and decision-making capabilities [7, 9]. Notably, RL-based models such as OpenAI-o1 and DeepSeek-R1 have demonstrated substantial gains in logical inference and iterative reasoning—achieved purely through reward-driven learning, without relying on explicit step-by-step supervision [19]. Within this paradigm, several studies have explored using RL to train policy models that can more effectively search for relevant information. Representative examples include Search-R1 [16], R1-Searcher [35], and ReSearch [18]. Notably, DeepResearcher [48] and WebThinker [23] introduce live interaction with commercial search engines such as Google, allowing models to train in an environment that closely mirrors real-world web search. Despite these advancements, integrating RL with real-world search scenarios presents several significant challenges: (1) Uncontrolled Document Quality: The quality of documents retrieved from live search engines is often unpredictable, introducing noise and instability into the training process. (2) Prohibitively High API Costs: RL training requires frequent rollouts, potentially involving hundreds of thousands of API calls, which incur substantial financial costs and severely limit scalability.
@@
To address these challenges, we propose ZeroSearch—a novel reinforcement learning framework that enables LLMs to learn search strategies without interacting with real search engines. Our key insight is that LLMs have acquired extensive world knowledge during large-scale pretraining and are capable of generating relevant documents given a search query [44]. The primary difference between a real search engine and a simulation LLM lies in the textual style of the returned content. However, with lightweight supervised fine-tuning, even relatively small LLMs can effectively simulate the behavior of real search engines. In addition to eliminating API costs, an important advantage of using LLMs for document generation is the ability to control document quality. Specifically, during supervised fine-tuning, documents that lead to correct or incorrect answers are distinguished through prompt design, enabling the simulation LLM to learn to generate either useful or noisy documents simply by adjusting a few words in the prompt. Building on this, we introduce a curriculum rollout mechanism during training, in which the quality of the generated documents is gradually degraded over time to simulate increasingly challenging retrieval scenarios. This allows the policy model to first learn basic output formats and task requirements before progressively adapting to more challenging and noisy retrieval scenarios. More importantly, ZeroSearch exhibits strong scalability: increasing the number of GPUs significantly accelerates the generation throughput of the simulation LLM, thereby enabling efficient large-scale rollout. Empirical results show that even a 3B LLM used as the simulated search engine can effectively incentivize the policy model’s capabilities to utilize a real search engine. When a 7B LLM is used as the simulator, the resulting policy model achieves performance comparable to that trained with a real search engine, while a 14B simulator even leads to better performance. ZeroSearch is compatible with both base and instruction-tuned models of various parameter sizes, removing the need for separate supervised warm-up stages. Moreover, it integrates seamlessly with widely used RL algorithms, including REINFORCE [38], Proximal Policy Optimization (PPO) [31], and Group Relative Policy Optimization (GRPO) [32, 7].
@@
  • We propose ZeroSearch, a novel reinforcement learning framework that incentivizes the capability of LLMs to use real search engines without interacting with them during training.
  • Through supervised fine-tuning, we transform the LLM into a retrieval module capable of generating both useful and noisy documents in response to a query. We further introduce a curriculum rollout mechanism to progressively elicit the model’s reasoning ability by exposing it to increasingly challenging retrieval scenarios.
@@ 2.2 Learning to Search through Reinforcement Learning
Recently, reinforcement learning (RL) has emerged as a promising paradigm for enhancing the reasoning capabilities of LLMs [7, 9]. Notable RL-based models such as OpenAI-o1 and DeepSeek-R1 have demonstrated remarkable capabilities in logical inference and iterative reasoning, purely driven by reward signals without explicit step-by-step supervision [19]. Several studies have also explored RL techniques specifically designed to train models for effective information retrieval. For instance, Search-R1 [16] employs reinforcement learning to autonomously generate multiple search queries during step-by-step reasoning. Similarly, R1-Searcher [35] proposes a two-stage, outcome-based RL method aimed at enhancing search capabilities. ReSearch [18] leverages RL to teach models to reason through searches, entirely without supervision on intermediate reasoning steps. However, these methods usually employ static, local textual corpora such as Wikipedia and fail to capture the complexities of real-world interaction. To bridge this gap, DeepResearcher [48] and WebThinker [23] introduce direct interaction with commercial search engines such as Google, allowing training environments that closely approximate real-world search scenarios. While achieving superior performance, these real-time retrieval methods face significant challenges, including unpredictable document quality, prohibitive high API costs that adversely affect system scalability. To address these limitations, we propose ZeroSearch, a method leveraging an LLM to simulate real-time search, effectively removing dependence on costly, rate-limited real search APIs. Through lightweight supervised fine-tuning, ZeroSearch allows explicit control over document quality and implements a curriculum rollout mechanism that enhances training stability and robustness.
@@ 3.3 Search Simulation Tuning
To bridge this gap, we propose a lightweight supervised fine-tuning (SFT) procedure. Specifically, we first collect interaction trajectories by prompting the LLM to engage with a real search engine in a multi-turn manner until a final answer is reached. From these trajectories, we extract query-document pairs and employ the LLM as the judge to independently assess whether each document contains sufficient information to answer the corresponding query. If the answer is affirmative, the document is labeled as a useful output; otherwise, it is labeled as a noisy output.
Then, we perform lightweight SFT to enhance the LLM’s ability to generate both useful and noisy outputs in response to queries. As shown in Table 2, the distinction between useful and noisy outputs can be effectively controlled by adjusting a few words in the prompt. Besides, we also incorporate the input question and its corresponding answer into the prompt to broaden the knowledge boundary of the simulation LLM. After fine-tuning, the simulation LLM is capable of generating both useful and noisy documents, enabling dynamic document quality control during rollout.
@@ 3.6 Training Algorithm
Our approach is compatible with a wide range of reinforcement learning algorithms, including REINFORCE [38], Proximal Policy Optimization (PPO) [31], and Group Relative Policy Optimization (GRPO) [32, 7], each offering distinct advantages for optimizing retrieval-augmented reasoning.
In ZeroSearch, the rollout sequence comprises both tokens generated by the policy model and document tokens returned by the simulation LLM. Applying the same optimization procedure uniformly across both types of tokens can lead to training instability, as the document tokens are externally generated and not directly controlled by the policy model.
To mitigate this, we introduce a loss masking mechanism for document tokens, ensuring that gradients are only computed with respect to the model’s own outputs. This strategy stabilizes the RL training process while preserving the effectiveness of retrieval-augmented generation.
@@ 4.3 Experimental Setup
We conduct experiments using three model families: Qwen-2.5-7B (Base/Instruct) and Qwen-2.5-3B (Base/Instruct) [41], as well as LLaMA-3.2-3B (Base/Instruct) [4]. To simulate real-world retrieval scenarios, we utilize Google Web Search via the SerpAPI [^1] https://serpapi.com/ as the external search engine. During evaluation, all methods use SerpAPI as the search engine to ensure a fair comparison. The number of retrieved documents is fixed at five across all methods to ensure a fair comparison.
@@
During experimentation, we deploy the simulation server on 4 H20 GPUs and conduct RL training on another 4 H20 GPUs. To train ZeroSearch, we experiment with three RL algorithms: REINFORCE, GRPO, and PPO. Unless otherwise specified, all RL-based methods use REINFORCE as the default training algorithm. The fine-tuned Qwen-2.5-14B-Instruct model serves as the default simulated search engine across all experiments. During inference, all models interact with the real web environment via Google Web Search to ensure a fair comparison. For further implementation details, including complete hyperparameter configurations, please refer to Appendix F.
@@ 4.4 Performance
R1-base 27.60 47.40 27.40 21.00 29.20 9.80 27.78 27.17
R1-instruct 27.00 45.80 24.20 21.60 27.80 8.40 25.00 25.69
Search-R1-base 43.40 61.40 54.60 31.20 37.20 18.20 30.56 39.51
Search-R1-inst 42.40 63.40 51.60 32.80 33.20 17.40 26.39 38.17
@@
R1-base 14.20 34.80 20.80 19.60 28.40 6.40 5.56 18.54
R1-instruct 19.80 33.00 19.40 19.40 26.40 4.40 11.11 19.07
Search-R1-base 40.60 60.00 44.20 29.20 32.00 11.20 12.50 32.81
Search-R1-inst 35.80 55.80 26.00 33.20 26.00 7.60 12.50 28.13
@@
R1-base 28.40 44.20 30.00 22.80 28.40 7.00 11.11 24.56
R1-instruct 35.00 52.20 27.60 21.60 17.80 11.40 20.83 26.63
Search-R1-base 41.20 60.00 44.00 29.60 31.60 13.60 19.44 34.21
Search-R1-inst 37.60 53.60 44.20 21.00 20.40 8.80 27.78 30.48
@@ 5.1 Compare ZeroSearch with Real Search Engine
We compare the reward curves of ZeroSearch and Search-R1 (using a real search engine) on Qwen-2.5-3B, as shown in Figures 2(a) and 2(b). Several key observations can be made:
@@
ZeroSearch achieves a more pronounced reward improvement. As shown in Figure 2(b), ZeroSearch initially lags behind Search-R1 but eventually surpasses it with less fluctuation, thanks to the curriculum rollout mechanism that helps the model gradually master search tool usage.
@@
(a) Qwen-2.5-3B-Base
(b) Qwen-2.5-3B-Inst
@@
Figure 2: (a-b): Reward curve comparison between ZeroSearch and Search-R1 using Qwen-2.5-3B. (c): The reward curve and interaction turns during training of LLaMA-3.2-3B-Base.
@@ 5.2 Choice of Simulation LLMs
SFT-7B 40.40 60.40 45.40 31.20 30.80 12.60 13.89 33.53
SFT-14B 43.00 61.60 41.40 33.80 34.60 13.00 13.89 34.47
[…]
Table 4: Performance of different simulated search engines with Qwen-2.5-3B-Base as the policy model. The Base Model refers to performance without any RL training.
In this section, we investigate how different simulation engine configurations affect the performance of the policy model, including prompt-based and fine-tuned LLMs ranging from 3B to 14B parameters. Based on the results in Table 4, we make the following observations:
@@
Second, the fine-tuned simulation engines significantly outperform prompt-based simulation engines. Although prompt-based methods are explicitly guided to mimic the response style of a real search engine, a substantial distribution gap remains, leading to inferior performance.
Third, performance improves consistently with increasing model size. Larger simulation LLMs not only exhibit stronger simulation capabilities but also more accurately distinguish between useful and noisy documents, thereby enabling more effective curriculum rollout during training.
@@ 5.3 Interaction Turns Study
In this section, we analyze the training dynamics of ZeroSearch by examining both the reward curve and the number of interaction turns throughout the training, using the LLaMA-3.2-3B-Base as the policy model. The results are shown in Figure 2(c).
@@
As training progresses, both the number of interaction turns and the reward curve increase sharply and then stabilize. This is primarily because the policy model becomes capable of effectively retrieving relevant documents and ultimately achieving correct answers, resulting in higher rewards. Notably, although the reward appears stable in the later stages of training, the underlying task difficulty continues to rise due to the curriculum rollout mechanism. Therefore, the policy must continuously refine its strategy and improve its reasoning ability to maintain the high reward score.
5.4 Different RL algorithms: REINFORCE vs. GRPO vs. PPO
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
REINFORCE 43.00 61.60 41.40 33.80 34.60 13.00 13.89 34.47
GRPO 39.40 55.80 39.00 29.80 32.20 12.40 23.61 33.17
PPO 38.20 58.60 40.40 27.20 33.80 13.80 16.67 32.67
Table 5: Performance of ZeroSearch under different RL algorithms. We compare REINFORCE, GRPO and PPO using Qwen-2.5-3B-Base as the policy model.
In this section, we evaluate the performance of three widely used RL training algorithms, REINFORCE, GRPO and PPO, within the ZeroSearch framework, using the Qwen-2.5-3B-Base as the policy model. The results of this comparison are presented in Table 5.
We observe that all three algorithms effectively enhance the model’s ability to perform search within our framework, demonstrating the strong generalization ability of ZeroSearch. Among them, REINFORCE achieves the best performance, consistent with the observations in Search-R1 [16], further highlighting its advantages in training stability. It is also worth noting that both REINFORCE and GRPO involve repeated rollouts, which incur higher API costs when interacting with a real search engine. This further underscores the practicality of our simulated search setup.
5.5 Curriculum Rollout Study